OpenStudy (anonymous):

Write the equation of the line that passes through the points (3, 6) and (4, 10) using function notation.

OpenStudy (anonymous):

please help!!!

OpenStudy (whovianchick):

slope=y1-y2/x1-x2 m=10-6/4-3 m=4 Point slope form: y-10=4(x-4) y-10=4x-16 y=4x-26 f(x)=4x-26

OpenStudy (anonymous):

thank you i have 5 minutes left

OpenStudy (whovianchick):

no prob

OpenStudy (anonymous):

wait @whovianchick

OpenStudy (anonymous):

that has 2 be wrong

OpenStudy (anonymous):

the answers are: f(x) = 4x − 6 f(x) = x + 4 y = x + 4 y = 4x − 6

OpenStudy (whovianchick):

yeah you're right one sec

OpenStudy (whovianchick):

It's A, it's not -26 it's -6 my bad

OpenStudy (anonymous):

the slope is 5.5 btw

OpenStudy (anonymous):

u got that wrong

OpenStudy (whovianchick):

no the slope is 4, look at the answers. None of them say y=5.5x

OpenStudy (whovianchick):

The slope is the number in front of x

OpenStudy (anonymous):

well yeah my bad

OpenStudy (anonymous):

your right

OpenStudy (whovianchick):

It's A, I plugged in the points. Super Sorry that I messed up b (Equation is y=mx+b where m is the slope and b is the y-intercept)

OpenStudy (anonymous):

ok thx